The line Ax + By =8 passes through the points (2,1) and (4,-2). Find the values of A & B. Show work please.
Found 2 solutions by solver91311, stanbon:
Answer by solver91311(12126)



You need to use the Two-Point form of the equation of a straight line because you are given two ordered pairs to define your line.



Where and are your given points.

It really doesn't matter which point you select as #1 and #2, so I just selected: and giving us:



Now all you have to do is the arithmetic and algebraic manipulation to put the equation into the requested form, namely:



John


Answer by stanbon(48568)

The line Ax + By =8 passes through the points (2,1) and (4,-2). Find the values of A & B. Show work please.
----------------------
Substitute those x/y values to get two equations.
2A + B = 8
4A -2B = 8
--------------------------
Multiply thru the 1st equation by 2 to get:
4A + 2B = 8
4A - 2B = 8
-----------------------------
Add the two equations to solve for A:
8A = 16
A = 2
--------
Substitute that value of A to solve for B.
2*2 + B = 8
B = 4
======================
Cheers,
San H.